# Multi-View Contrastive Network (MCNet) for Motor Imagery Classification

### v1

@@ Abstract
Objective: An electroencephalography (EEG)-based brain-computer interface (BCI) serves as a direct communication pathway between the human brain and an external device. While supervised learning has been extensively explored for motor imagery (MI) EEG classification, small data quantity has been a key factor limiting the performance of deep feature learning. Methods: This paper proposes a knowledge-driven time-space-frequency based multi-view contrastive network (MCNet) for MI EEG decoding in BCIs. MCNet integrates knowledge from the time, space, and frequency domains into the training process through data augmentations from multiple views, fostering more discriminative feature learning of the characteristics of EEG data. We introduce a cross-view contrasting module to learn from different augmented views and a cross-model contrasting module to enhance the consistency of features extracted between knowledge-guided and data-driven models. Results: The combination of EEG data augmentation strategies was systematically investigated for more informative supervised contrastive learning. Experiments on four public MI datasets and three different architectures demonstrated that MCNet outperformed 10 existing approaches. Significance: Our approach can significantly boost EEG classification performance beyond designated networks, showcasing the potential to enhance the feature learning process for better EEG decoding.
@@ I Introduction
Contrastive learning has shown great advances in extracting powerful representations of data, mostly in the form of self-supervised learning using unlabeled data via contrasting different augmented views [12, 13, 14]. In this work, we study supervised contrastive learning to fully unleash the potential of additional integration of prior expert knowledge in data-driven learning, as illustrated in Fig. 1. Our research shows that the incorporation of knowledge-guided data augmentations is paramount to enable neural networks to learn invariant transformations of EEG data. Specifically, we propose time-space-frequency based multi-view contrastive network (MCNet) to solve the current obstacles that hinder more discriminative EEG feature learning:
@@
    We propose MCNet, introducing a cross-view contrasting module to explore features of different augmented views, and a cross-model contrasting module to enhance the consistency of features extracted between knowledge-guided and data-driven models.
@@
    Empirical results demonstrate that MCNet consistently outperformed 10 baselines on three different architectures and with various data augmentation strategies on four public MI datasets.
The remainder of this paper is organized as follows: Section II reviews related works. Section III details the proposed MCNet approach. Section IV discusses the experimental results and provides analyses. Section V draws conclusions.
@@ III Methodology
Figure 2: Multi-view contrastive network (MCNet) for EEG-based BCIs.
@@ III-C Cross-Model Contrasting
III-D MCNet
Integrating cross-view contrasting, cross-model contrasting, and supervised classification modules, the overall loss function of MCNet is:
@@ III-D MCNet
The pseudo-code of MCNet is given in Algorithm 1.
Algorithm 1 Multi-View Contrastive Network (MCNet).
@@ IV Experiments and Results
This section presents the datasets, experiments and analyses. Code is available on GitHub [^1] https://github.com/wzwvv/MCNet.
@@ IV-B Algorithms
We compared MCNet with 10 approaches on three networks, including Baseline, data augmentation strategies, and contrastive learning approaches:
@@
Note that for SimCLR and InfoNCE, the number of branches under the proposed SCL framework is reduced to two, instead of three in MCNet. The two views are randomly selected each time from the data augmentation pool.
@@ IV-C Implementation
Implementation details. We performed leave-one-subject-out cross-validation on all datasets, as it is a more valuable and practical setting for EEG-based BCI applications. All experiments were repeated five times and average results were reported. For all datasets, the models were trained for 100 epochs using Adam optimizer with learning rate $10^{-3}$. $\tau$ was set to 0.2 as in [41]. $\lambda$ and $\gamma$ were set to 0.1 for all datasets. The batch size was 32 for Baseline, 64 for data augmentation approaches, and 256 for SimCLR, InfoNCE and MCNet based on the size of training data. Note that for MCNet, three distinct data augmentation strategies are applied: Scale (time domain), Shift (frequency domain), and HS (space domain).
@@ IV-D Main Results
    The results of SimCLR, InfoNCE and the proposed MCNet were better than other 7 baselines under augmentation techniques, demonstrating the effectiveness of the proposed SCL framework.
@@
    Our proposed MCNet consistently outperformed 10 baseline methods across all backbone networks. In comparison with SimCLR, the most similar approach, MCNet ensures that the captured features originate from three distinct domains, while SimCLR only considers two views that might come from the same domain. This suggests that MCNet integrates more valuable information and learns more discriminative representations.
@@ IV-E Combinations of Multiple Data Augmentations
To explore the influence of the combination strategy of data augmentation, we evaluated the performance of all conceivable pairwise combinations of the seven data augmentation techniques in the proposed SCL framework. The results in Fig. 4 show that the effectiveness of data augmentation is more pronounced when utilizing combinations rather than relying on individual augmentation methods. For Zhou2016, the optimal combination achieved an accuracy of 87.01%, which is lower than the performance of MCNet when utilizing all three views simultaneously. Similar results can be observed for other datasets. It indicates the necessity of applying multiple forms of transformations, and also verifies the effectiveness of our proposed approach beyond bare augmentations.
@@ IV-G Visualizations
Fig. 7 shows the $t$-SNE [42] visualizations of the Baseline (EEGNet) feature distributions and MCNet feature distributions under different augmentation techniques in Blankertz2007 and BNCI2014002 datasets. Observe that the feature distributions of the proposed MCNet are structurally tighter and can be classified more easily than the feature distributions extracted from EEGNet. It suggests the importance of considering all three views together and MCNet can effectively minimize their gaps to learn the generalized representations.
Figure 7: $t$-SNE visualization of (a) EEGNet features and (b) MCNet features extracted from Subject S1-S7 of Blankertz2007; (c) EEGNet features and (d) MCNet features extracted from Subject S1-S14 of BNCI2014002. Different colors represent trials transformed through different augmentation approaches, including CR (spatial domain), Scale (time domain), and Shift (frequency domain). The dots and crosses represent trials of different categories.
@@ V Conclusion
Acquiring labeled calibration data from new subjects is a labor-intensive and time-consuming task. The limited quantity of available data has constrained the deep feature learning capabilities of EEG signals, hindering the widespread development of EEG-based BCIs. This paper introduces a novel MCNet framework to address the common problem of insufficient data in EEG-based BCIs. MCNet strategically incorporates knowledge from the time, space, and frequency domains into the training process, thereby augmenting the discriminative features extracted from EEG signals. The proposed framework consists of an end-to-end network, featuring a cross-view contrasting module for exploring discriminative features across distinct augmented views, a cross-model contrasting module to maximize feature consistency across diverse models, and a supervised classification module.
Through experiments conducted on four publicly available MI datasets, MCNet demonstrated superior performance compared to 10 existing approaches. Moreover, it showcased flexibility in design by seamlessly incorporating various data augmentation strategies and proving effective on different backbone networks. These findings underscore the potential of MCNet in significantly enhancing the effectiveness and practicality of future BCI systems.

### v2

# Multi-View Contrastive Network (MVCNet) for Motor Imagery Classification

@@ Abstract
Objective: An electroencephalography (EEG)-based brain-computer interface (BCI) serves as a direct communication pathway between the human brain and an external device. While supervised learning has been extensively explored for motor imagery (MI) EEG classification, small data quantity has been a key factor limiting the performance of deep feature learning. Methods: This paper proposes a knowledge-driven time-space-frequency based multi-view contrastive network (MVCNet) for MI EEG decoding in BCIs. MVCNet integrates knowledge from the time, space, and frequency domains into the training process through data augmentations from multiple views, fostering more discriminative feature learning of the characteristics of EEG data. We introduce a cross-view contrasting module to learn from different augmented views and a cross-model contrasting module to enhance the consistency of features extracted between knowledge-guided and data-driven models. Results: The combination of EEG data augmentation strategies was systematically investigated for more informative supervised contrastive learning. Experiments on four public MI datasets and three different architectures demonstrated that MVCNet outperformed 10 existing approaches. Significance: Our approach can significantly boost EEG classification performance beyond designated networks, showcasing the potential to enhance the feature learning process for better EEG decoding.
@@ I Introduction
Contrastive learning has shown great advances in extracting powerful representations of data, mostly in the form of self-supervised learning using unlabeled data via contrasting different augmented views [12, 13, 14]. In this work, we study supervised contrastive learning to fully unleash the potential of additional integration of prior expert knowledge in data-driven learning, as illustrated in Fig. 1. Our research shows that the incorporation of knowledge-guided data augmentations is paramount to enable neural networks to learn invariant transformations of EEG data. Specifically, we propose time-space-frequency based multi-view contrastive network (MVCNet) to solve the current obstacles that hinder more discriminative EEG feature learning:
@@
    We propose MVCNet, introducing a cross-view contrasting module to explore features of different augmented views, and a cross-model contrasting module to enhance the consistency of features extracted between knowledge-guided and data-driven models.
@@
    Empirical results demonstrate that MVCNet consistently outperformed 10 baselines on three different architectures and with various data augmentation strategies on four public MI datasets.
The remainder of this paper is organized as follows: Section II reviews related works. Section III details the proposed MVCNet approach. Section IV discusses the experimental results and provides analyses. Section V draws conclusions.
@@ III Methodology
Figure 2: Multi-view contrastive network (MVCNet) for EEG-based BCIs.
@@ III-C Cross-Model Contrasting
III-D MVCNet
Integrating cross-view contrasting, cross-model contrasting, and supervised classification modules, the overall loss function of MVCNet is:
@@ III-D MVCNet
The pseudo-code of MVCNet is given in Algorithm 1.
Algorithm 1 Multi-View Contrastive Network (MVCNet).
@@ IV Experiments and Results
This section presents the datasets, experiments and analyses. Code is available on GitHub [^1] https://github.com/wzwvv/MVCNet.
@@ IV-B Algorithms
We compared MVCNet with 10 approaches on three networks, including Baseline, data augmentation strategies, and contrastive learning approaches:
@@
Note that for SimCLR and InfoNCE, the number of branches under the proposed SCL framework is reduced to two, instead of three in MVCNet. The two views are randomly selected each time from the data augmentation pool.
@@ IV-C Implementation
Implementation details. We performed leave-one-subject-out cross-validation on all datasets, as it is a more valuable and practical setting for EEG-based BCI applications. All experiments were repeated five times and average results were reported. For all datasets, the models were trained for 100 epochs using Adam optimizer with learning rate $10^{-3}$. $\tau$ was set to 0.2 as in [41]. $\lambda$ and $\gamma$ were set to 0.1 for all datasets. The batch size was 32 for Baseline, 64 for data augmentation approaches, and 256 for SimCLR, InfoNCE and MVCNet based on the size of training data. Note that for MVCNet, three distinct data augmentation strategies are applied: Scale (time domain), Shift (frequency domain), and HS (space domain).
@@ IV-D Main Results
    The results of SimCLR, InfoNCE and the proposed MVCNet were better than other 7 baselines under augmentation techniques, demonstrating the effectiveness of the proposed SCL framework.
@@
    Our proposed MVCNet consistently outperformed 10 baseline methods across all backbone networks. In comparison with SimCLR, the most similar approach, MVCNet ensures that the captured features originate from three distinct domains, while SimCLR only considers two views that might come from the same domain. This suggests that MVCNet integrates more valuable information and learns more discriminative representations.
@@ IV-E Combinations of Multiple Data Augmentations
To explore the influence of the combination strategy of data augmentation, we evaluated the performance of all conceivable pairwise combinations of the seven data augmentation techniques in the proposed SCL framework. The results in Fig. 4 show that the effectiveness of data augmentation is more pronounced when utilizing combinations rather than relying on individual augmentation methods. For Zhou2016, the optimal combination achieved an accuracy of 87.01%, which is lower than the performance of MVCNet when utilizing all three views simultaneously. Similar results can be observed for other datasets. It indicates the necessity of applying multiple forms of transformations, and also verifies the effectiveness of our proposed approach beyond bare augmentations.
@@ IV-G Visualizations
Fig. 7 shows the $t$-SNE [42] visualizations of the Baseline (EEGNet) feature distributions and MVCNet feature distributions under different augmentation techniques in Blankertz2007 and BNCI2014002 datasets. Observe that the feature distributions of the proposed MVCNet are structurally tighter and can be classified more easily than the feature distributions extracted from EEGNet. It suggests the importance of considering all three views together and MVCNet can effectively minimize their gaps to learn the generalized representations.
Figure 7: $t$-SNE visualization of (a) EEGNet features and (b) MVCNet features extracted from Subject S1-S7 of Blankertz2007; (c) EEGNet features and (d) MVCNet features extracted from Subject S1-S14 of BNCI2014002. Different colors represent trials transformed through different augmentation approaches, including CR (spatial domain), Scale (time domain), and Shift (frequency domain). The dots and crosses represent trials of different categories.
@@ V Conclusion
Acquiring labeled calibration data from new subjects is a labor-intensive and time-consuming task. The limited quantity of available data has constrained the deep feature learning capabilities of EEG signals, hindering the widespread development of EEG-based BCIs. This paper introduces a novel MVCNet framework to address the common problem of insufficient data in EEG-based BCIs. MVCNet strategically incorporates knowledge from the time, space, and frequency domains into the training process, thereby augmenting the discriminative features extracted from EEG signals. The proposed framework consists of an end-to-end network, featuring a cross-view contrasting module for exploring discriminative features across distinct augmented views, a cross-model contrasting module to maximize feature consistency across diverse models, and a supervised classification module.
Through experiments conducted on four publicly available MI datasets, MVCNet demonstrated superior performance compared to 10 existing approaches. Moreover, it showcased flexibility in design by seamlessly incorporating various data augmentation strategies and proving effective on different backbone networks. These findings underscore the potential of MVCNet in significantly enhancing the effectiveness and practicality of future BCI systems.